# The nature of mathematical models

### v1

@@ 3.1 Preliminaries
From what has been presented in the Introduction, modeling appears to be the act of hypothesizing mathematical relationships among idealized quantities, some of which may be observed (i.e. numerically measured in a certain number of cases). To make this idea precise, a few relevant, very standard results are listed in the following. Proofs can be found in any textbook of functional analysis or mathematical statistics [15, 16].
@@
Recall that a $n$-dimensional random variable is a measurable map from the probability space $(\Omega,\mathcal{F},P)$ to the measurable space $(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}))$, where $\mathcal{B}(\mathbb{R}^{n})$ is the Borel $\sigma$-algebra of $\mathbb{R}^{n}$.
The Doob-Dynkin lemma for random variables (also see Fig.3)
@@
Let $(\Omega,\mathcal{F},P)$ be a probability space, $(\mathbb{R}^{n},\mathcal{B}_{n})$ and $(\mathbb{R}^{m},\mathcal{B}_{m})$ measurable spaces, $U:\Omega\rightarrow\mathbb{R}^{n}$ a random variable. Then, a function $X:\Omega\rightarrow\mathbb{R}^{m}$ is $\sigma(U)$-$\mathcal{B}_{m}$-measurable iff $\,\,\exists\,x:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, $g$ $\mathcal{B}_{n}$-$\mathcal{B}_{m}$-measurable, such that $X=x\circ U=x(U)$. Consequently, $X$ is also a random variable.
@@
Figure 3: By the Doob-Dynkin lemma this diagram commutes
The space $\mathcal{L}^{p}(\Omega,\mathcal{F},\mathbb{P})$
[…]
We define the equivalence class of a measurable function $X$ under the equivalence relationship =$\overset{a.e.}{=}$ as $[X]\in\mathcal{L}^{p}/\overset{a.e.}{=}$ as $[X]=\{Y\in\mathcal{L}^{p}:Y\overset{a.e.}{=}X\}$.
[…]
By an abuse of notation we often call $\mathcal{L}^{p}$ the quotient space $\mathcal{L}^{p}/\overset{a.e.}{=}$ and say that $X\in\mathcal{L}^{p}$ when we really mean $[X]\in\mathcal{L}^{p}/\overset{a.e.}{=}$.
It can finally be shown that $\mathcal{L}^{p}(\Omega,\mathcal{F},P))$, $p\geq 1$ is a Banach space (i.e complete and normed).
The Hilbert space of finite-variance random variables, $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathcal{F},\mathbb{P})$
[…]
Notice that $\mathcal{L}^{p}(\Omega,\mathcal{F},\mathbb{P})$ is not a Hilbert space for $p\neq 2$: we will work strictly in $\mathcal{L}^{2}(\Omega,\mathcal{F},P)$.
@@
[…]
@@
For every point $x$ in a Hilbert space $\mathcal{H}$ and every nonempty closed convex subset $C\subseteq H$ there exists a unique vector $y\in C$ for which $\lVert x-y\rVert$ is equal to $\delta:=\underset{c\in C}{\inf}\|x-c\|$
@@
Let $\mathcal{H}_{1}\subset\mathcal{H}$ be a subspace of a Hilbert space $\mathcal{H}$. Then, $\forall x\in\mathcal{H},\ \exists$ a unique vector $\mathcal{P}x\in\mathcal{H}_{1}$ such that $\forall z\in\mathcal{H}_{1},\ \langle x-\mathcal{P}x,z\rangle=0$. $\mathcal{P}x$ is called the projection of $x$ onto $\mathcal{H}_{1}$. In addition, $\mathcal{P}x$ is the element of minimum distance from $\mathcal{H}_{1}$
@@
[…]
@@
If $C$ is a closed vector subspace of a Hilbert space, then it is a Hilbert subspace. In addition, every finite dimensional vector subspace of an Hilbert space is closed
@@ 3.2 The nature of mathematical models: definitions
We have now the motivation and have available all the necessary structures for a formal treatment of what mathematical models for the applied sciences are. We refer to the above definition of $\mathcal{H}$.
@@
Given a random variable of interest or objective random variable, $\boldsymbol{X}^{o}$, possibly supposed to be related to some random variable $\boldsymbol{U}$, a model of $\boldsymbol{X}^{o}$ is an operator $\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{U}):\mathcal{H}\times\mathbb{R}^{q}\rightarrow\mathcal{H}$, whose arguments are some parameter $\boldsymbol{\theta}\in\mathbb{R}^{q}$ together with the predictor random variable(s) $\boldsymbol{U}$ (if any), and whose value is another random variable $\boldsymbol{X}=\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{U})$.
@@
Notice that the above is a slight abuse of notation: since we are typically interested in the relationship between the parameter $\boldsymbol{\theta}$ and the value $\boldsymbol{X}$ of the model $\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{U})$, having for all intents and purposes fixed $\boldsymbol{U}$, we just consider $\boldsymbol{X}$ as a function of $\boldsymbol{\theta}$. To keep the two concepts distinct however, we use the two different names "model" and "model map".
@@
Having fixed $\boldsymbol{U}$ and given sufficient regularity of the model, the set of values =$\mathcal{M}=\{\boldsymbol{X}(\boldsymbol{\theta}),\boldsymbol{\theta}\in\boldsymbol{\Theta}\}$, indexed by the parameter, is the model manifold, $\mathcal{M}\subset\mathcal{H}$
@@ 3.3 Linear and nonlinear models
If the hypothesized model is linear, e.g. $\boldsymbol{X}(\boldsymbol{\theta})=\boldsymbol{U}\,\boldsymbol{\theta}$ , then it is easy to see that $\mathcal{M}$ is a vector space with elements in $\mathcal{H}$, hence a subspace of the Hilbert space $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathcal{F},\mathbb{P})$.
@@
As a subspace of a Hilbert space, $\mathcal{M}$ in this case inherits its Hilbert space structure (norm, inner product), and a projection operator is guaranteed. Consequently, the point of projection $\hat{\boldsymbol{X}}=\mathcal{P}_{{\mathcal{M}}}(\boldsymbol{X}^{o})$ is the one of minimal distance from $\boldsymbol{X}^{o}$ (in the norm of $\mathcal{H}$), is unique and    ($\boldsymbol{X}^{o}-\hat{\boldsymbol{X}}$)   is orthogonal to all elements of $\mathcal{M}$.
@@
If on the other hand the model is nonlinear, e.g. $\boldsymbol{X}(\theta)=\sin(\theta\boldsymbol{U})$  , then $\mathcal{M}$ is not in general convex, much less a subspace of $\mathcal{H}$. In our example in fact :=$\mathcal{M}:=\{\boldsymbol{X}(\theta)\}$ is not convex since for $\theta^{1},\theta^{2}\in\boldsymbol{\Theta},\theta^{1}\neq\theta^{2}$ we have $X(\theta^{1})+X(\theta^{2})\notin\mathcal{M}$. Hence $\mathcal{M}$ is not a vector space, not a subspace of the Hilbert space $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathcal{F},\mathbb{P})$, no projection operator $\mathcal{P}_{{\mathcal{M}}}$ can be defined and no unique “best" estimate $\hat{\boldsymbol{X}}$ can be obtained.
@@
However, assume $\exists$ open $\,\mathcal{U}\subset\mathcal{M}$ and $\hat{\boldsymbol{X}}=\boldsymbol{X}(\hat{\boldsymbol{\theta}})\in\mathcal{U}$ such that $\forall\boldsymbol{X}\in\mathcal{U}$, $\boldsymbol{X}\neq\hat{\boldsymbol{X}}$ it is $\mathrm{d}(X^{o},\hat{X})<\mathrm{d}(X^{o},X)$. Hence, we assume we have an isolated point $\hat{X}\in\mathcal{M}$ locally closest to $X^{o}$. Under this hypothesis, that a random variable $\hat{\boldsymbol{X}}$ exists, of minimal distance from $\boldsymbol{X}^{o}$ within a neighborhood (in the topology on $\mathcal{H}$ induced by the norm), then we could consider working on a local linear approximation to the manifold $\mathcal{M}$ at $\hat{\boldsymbol{X}}$, that is we could consider constructing the tangent space to the manifold $\mathcal{M}$ at that locally optimal point, $T_{\hat{\boldsymbol{X}}}\mathcal{M}$. To do so, we need to use the concept of the tangent space to a generic topological manifold at one of its points. Any good introductory text in differential geometry (e.g. [17, 18, 19]) will provide the relevant definitions and proofs. Notice, again, that in general $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ is not a subspace of $\mathcal{H}$, but rather an affine space. However, the Hilbert projection theorem extends naturally to affine subspaces by translation, see Appendix 5.1 for details: here we will directly deal with the local approximation $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ to $\mathcal{M}$ at $\hat{\boldsymbol{X}}$, calling, with a slight abuse of language, the tangent space $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ to $\mathcal{M}$ (which is a bona-fide vector space in its own right) a “subspace” of $\mathcal{H}$, see Fig.LABEL:fig:NonLinearHilbertReparametrized.
@@
We can build $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ by considering the set of velocities, i.e. of differential operators, which map functions in $C^{\infty}(\mathcal{M})$ , $f:\mathcal{M}\rightarrow\mathbb{R}$ to their directional derivatives in the direction of curves $\gamma:\mathbb{R}\rightarrow\mathcal{M}$, $T_{\hat{\boldsymbol{X}}}\mathcal{M}=\{\vartheta_{\gamma,\hat{\boldsymbol{X}}}$ such that $\,\hat{\boldsymbol{X}}=\gamma(\lambda)\in\mathcal{U}\subset\mathcal{M}$ and $\forall f\in C^{\infty}(\mathcal{M})$ it is $\,\vartheta_{\gamma,\hat{\boldsymbol{X}}}=(f\circ\gamma)^{\prime}|_{\lambda}\}$. We can in general define a chart-induced basis of the tangent space: in the present case we will compute the basis for $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ induced by the canonical chart $\boldsymbol{X}^{-1}$:
$\mathcal{U}$$\mathbb{R}^{q}$$\boldsymbol{X}^{-1}$$\boldsymbol{X}$
@@
Clearly we must require sufficient regularity in the model map $\boldsymbol{X}$ to do this, such as being continuously differentiable with a continuously differentiable inverse. Notice that $\boldsymbol{X}^{-1}$ is a chart, indeed a global chart, mapping the whole of $\mathcal{M}$ to $\mathbb{R}^{q}$, since $\mathcal{M}$ by definition is generated by $\boldsymbol{X}(\boldsymbol{\Theta})$, with $\boldsymbol{\Theta}\in\mathbb{R}^{q}$. We can also say, therefore that:
@@
It is now necessary to clarify the correspondence between the (differential) operators in $T_{\hat{\boldsymbol{X}}}(\mathcal{M})$ and the random variables that ought to belong to the (affine) subspace of $\mathcal{H}$ it defines.
We can do this by using the “canonical" chart from the manifold to $\mathbb{R}^{q}$ induced by the parametrization of the model. As is detailed in Appendix LABEL:sec:AffineTangentToM We can express the directional derivative of a function $f$ in the direction of a curve $\gamma$ at a point $p=\gamma(\lambda)$, with respect to a specific chart map $\mu$, as the dot product of the components of the derivative of the curve representation under the chart $\gamma_{\mu}$ at $\lambda$, times the application of the operator $\frac{\partial}{\partial\mu}|_{p}=(\cdot\circ\mu^{-1})|_{p_{\mu}}$ to the function $f$.
For our specific application we can use as basis of the tangent space the basis induced by the chart map $(\boldsymbol{X}^{-1})$:
$\{(\cdot\circ((\boldsymbol{X}^{-1})^{-1})^{i})^{\prime}|_{\hat{\boldsymbol{X}}}\}_{i=1,...,q}=\{(\cdot\circ\boldsymbol{X}^{i})^{\prime}|_{\hat{\boldsymbol{X}}}\}_{i=1,...,q}\,\,.$

### v2

@@ 3.1 Preliminaries
From what has been presented in the Introduction, modeling appears to be the act of hypothesizing mathematical relationships among idealized quantities, some of which may be observed (i.e. numerically measured in a certain number of cases). To make this idea precise, a few relevant, very standard results are listed in the following. Proofs can be found in any textbook of functional analysis or mathematical statistics [Folland2013, Rohatgi1976].
@@
Recall that an $n$-dimensional real random variable $U:\Omega\rightarrow\mathbb{R}^{n}$ is a measurable map from the probability space $(\Omega,\mathcal{F},P)$ to the measurable space $(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}))$, where $\mathcal{B}(\mathbb{R}^{n})$ is the Borel $\sigma$-algebra of $\mathbb{R}^{n}$. Recall also that, given a probability space $(\Omega,\mathcal{F},P)$ and a random variable $U:\Omega\rightarrow\mathbb{R}^{n}$, the $\sigma$-algebra generated by the r.v. $U$, $\sigma(U)$, is the smallest $\sigma$-algebra containing all sets $U^{-1}(A),A\in\mathcal{B}(\mathbb{R}^{n})$.
The Doob-Dynkin lemma for random variables (also see Fig.3, page 3)
@@
Let $(\Omega,\mathcal{F},P)$ be a probability space, $(\mathbb{R}^{n},\mathcal{B}_{n})$ and $(\mathbb{R}^{m},\mathcal{B}_{m})$ measurable spaces, $U:\Omega\rightarrow\mathbb{R}^{n}$ a random variable. Then a function $X:\Omega\rightarrow\mathbb{R}^{m}$ is $\sigma(U)$-$\mathcal{B}_{m}$-measurable iff $\exists\,x:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, $x$ $\mathcal{B}_{n}$-$\mathcal{B}_{m}$-measurable, such that $X=x\circ U=x(U)$. In this case, consequently, $X$ is also a random variable.
@@
Figure 3: By the Doob-Dynkin lemma this diagram commutes for measurable functions
The spaces $\mathcal{L}^{p}(\Omega,\mathcal{F},P)$ and $L^{p}(\Omega,\mathcal{F},P)$
The set of measurable functions $X$ on $\Omega$ such that $(\int_{\Omega}|X|^{p}dP)^{\frac{1}{p}}<\infty$ is denoted by $\mathcal{L}^{p}(\Omega,\mathcal{F},P)$. It can be easily seen that $\mathcal{L}^{p}$ is a vector space.
Let $Y,X\in\mathcal{L}^{p}(\Omega,\mathcal{F},P)$. If $Y$ and $X$ differ only on a set of $P$-measure zero (i.e. $Y\overset{Pa.e.}{=}X$) then $(\int_{\Omega}|Y-X|^{p}dP)^{\frac{1}{p}}=0$. This leads naturally to consider such measurable functions as equivalent if they are equal almost everywhere. We can thus quotient $\mathcal{L}^{p}$ by the equivalence relation $\overset{Pa.e.}{=}$: the quotient space $L^{p}=\mathcal{L}^{p}/\overset{Pa.e.}{=}$ is defined to be the set of all equivalence classes under the relation $\overset{Pa.e.}{=}$.
We define the equivalence class of a measurable function $X$ under the equivalence relationship $\overset{Pa.e.}{=}$ as $X]\in L^{p}$, $X]=\{Y\in\mathcal{L}^{p}:Y\overset{Pa.e.}{=}X\}$.
We recall that a seminorm $s$ is a function from a vector space $\mathcal{V}$ over the reals to $\mathbb{R}$$^{+}$, $s:\mathcal{V}\to{\mathbb{R}}^{+}$, such that $\forall a,b\in\mathcal{V},\alpha\in{\mathbb{R}}$:
  1. 1.
    $s(a+b)\leqslant s(a)+s(b)$    (triangle inequality)
  2. 2.
    $s(\alpha a)=|\alpha|a$    (absolute homogeneity)
Notice that from property $(2)$ non-negativity follows: $s(a)\geqslant 0$ and $s(\boldsymbol{0})=0$.
If furthermore the function $s$ separates points, i.e.
  1. 3.
    $s(a)=0\Rightarrow a=\boldsymbol{0}$
then $s$ is a norm, which we typically indicate with double bars, $||\cdot||$.
While $\mathcal{L}^{p}$ is only a seminormed space, $\|[X]\|_{p}:=\|X\|_{p}=(\int_{\Omega}|X|^{p}d\mu)^{\frac{1}{p}}$ for any $X\in[X]$   is a norm in the quotient space $L^{p}=\mathcal{L}^{p}/\overset{Pa.e.}{=}$.
By an abuse of notation we will say in the following $X\in L^{p}$ when we really mean $X]\in L^{p}$.
It can finally be shown that the normed space $L^{p}(\Omega,\mathcal{F},P))$, $p\geq 1$, is complete, hence it is a Banach space.
The Hilbert space of finite-variance random variables, $\mathcal{H}=L^{2}(\Omega,\mathcal{F},P)$
Suppose that in the above construction we take $p=2$. Clearly, the set of measurable functions $X$ on $\Omega$ such that $(\int_{\Omega}|X|^{2}dP)^{\frac{1}{2}}<\infty$ coincides with the set $\{X:\int_{\Omega}X^{2}dP<\infty\}$, which is the set of all square-integrable measurable functions on $\Omega$ or the set of all finite-variance random variables on $(\Omega,\mathcal{F},P)$ .
If we endow $L^{2}(\Omega,\mathcal{F},P)$ with the inner product $\langle X,Y\rangle=\int_{\Omega}XY\,dP$, which is obviously compatible with the norm, we make it into an inner product space. Since, as noted above, it is a Banach space, it is also a Hilbert space.
We denote as $\mathcal{H}$, $\mathcal{H}=L^{2}(\Omega,\mathcal{F},P)$, this Hilbert space of random variables with finite variance, with $\langle X,Y\rangle=E(XY)=\int_{\Omega}XY\,dP$.
Notice that $L^{p}(\Omega,\mathcal{F},P)$ is not a Hilbert space for $p\neq 2$: we will thus work strictly in $L^{2}(\Omega,\mathcal{F},P)$.
@@
The distance between two elements of an Hilbert space $x,y\in\mathcal{X}$ is $\mathrm{d}(x,y)=\|x-y\|$, where $\|\cdot\|$ is the canonical norm
@@
For every point $x$ in a Hilbert space $\mathcal{H}$ and every nonempty closed convex subset $C\subseteq H$ there exists a unique vector $y\in C$ for which $\lVert x-y\rVert$ is equal to $\mathrm{d}(x,C):=\underset{c\in C}{\inf}\|x-c\|$
@@
Let $\mathcal{H}_{1}\subset\mathcal{H}$ be a subspace of a Hilbert space $\mathcal{H}$. Then, $\forall x\in\mathcal{H},\ \exists$ a unique vector $\mathcal{P}x\in\mathcal{H}_{1}$ such that $\forall z\in\mathcal{H}_{1},\ \langle x-\mathcal{P}x,z\rangle=0$. $\mathcal{P}x$ is called the projection of $x$ onto $\mathcal{H}_{1}$. In addition, $\mathcal{P}x$ is the element of minimum distance from $\mathcal{H}_{1}$ to $x$.
@@
  • $\mathcal{P}$ is a bounded and self-adjoint operator ($\forall\,x,y\in\mathcal{H}\,\,\,\left<\mathcal{P}x,y\right>=\left<x,\mathcal{P}y\right$>)
@@
If $C$ is a closed vector subspace of a Hilbert space, then it is a Hilbert subspace. In addition, every finite dimensional vector subspace of an Hilbert space is closed.
@@ 3.2 The nature of mathematical models: definitions
We have now the motivation and have available all the necessary tools for a formal treatment of what mathematical models for the applied sciences are. We refer to the above definition of $\mathcal{H}$.
@@
Given a random variable of interest or objective random variable, $\boldsymbol{X}^{o}$, possibly supposed to be related to some random variable $\boldsymbol{U}$, a model of $\boldsymbol{X}^{o}$ is an operator $\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U}):\mathbb{R}^{q}\times\mathcal{H}\rightarrow\mathcal{H}$, whose arguments are some parameter $\boldsymbol{\theta}\in\mathbb{R}^{q}$ together with the predictor random variable(s) $\boldsymbol{U}$ (if any), and whose value is another random variable $\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U})$.
@@
Notice that the above is a slight abuse of notation: since we are typically interested in the relationship between the parameter $\boldsymbol{\theta}$ and the value $\boldsymbol{X}$ of the model $\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U})$, having for all intents and purposes fixed $\boldsymbol{U}$ we just consider $\boldsymbol{X}$ as a function of $\boldsymbol{\theta}$. To keep the two concepts distinct, however, we use the two different names "model" and "model map".
For clarity, we distinguish between the model map defined above (having values in $\mathcal{H}$) and the corresponding model function, of the same functional form but having values in ${\mathbb{R}}^{n}$:
@@
A model function is a map $\boldsymbol{x}(\boldsymbol{\theta}):\mathbb{R}^{q}\rightarrow{\mathbb{R}}^{n}$, whose argument is some parameter $\boldsymbol{\theta}\in\mathbb{R}^{q}$ (considering any possible predictors fixed) and whose value is a point in ${\mathbb{R}}^{n}$, $\boldsymbol{x}:{\mathbb{R}}^{q}\to{\mathbb{R}}^{n},\,\,\,=\boldsymbol{x}:\boldsymbol{\theta}\mapsto\boldsymbol{x}(\boldsymbol{\theta})$.
Definition.
Having fixed $\boldsymbol{U}$ and assuming sufficient regularity of the model, the set of values ${\mathcal{M}_{\boldsymbol{X}}}=\{\boldsymbol{X}(\boldsymbol{\theta}),\boldsymbol{\theta}\in\boldsymbol{\Theta}\}$, indexed by the parameter, is the model manifold, $\mathcal{M}_{\boldsymbol{X}}$$\subset\mathcal{H}$
@@ 3.3 Linear and nonlinear models
If the hypothesized model is linear, e.g. $\boldsymbol{X}(\boldsymbol{\theta})=\boldsymbol{U}\,\boldsymbol{\theta}$ , then it is easy to see that $\mathcal{M}_{\boldsymbol{X}}$} is a vector space with elements in $\mathcal{H}$, hence a subspace of the Hilbert space $\mathcal{H}=L^{2}(\Omega,\mathcal{F},P)$.
@@
As a subspace of a Hilbert space, $\mathcal{M}_{\boldsymbol{X}}$} in this case inherits its Hilbert space structure (norm, inner product), and a projection operator is guaranteed. Consequently, the point of projection $\hat{\boldsymbol{X}}=\mathcal{P}_{{\mathcal{M}}_{\boldsymbol{X}}}(\boldsymbol{X}^{o})$ is the one of minimal distance from $\boldsymbol{X}^{o}$ (in the norm of $\mathcal{H}$), is unique and    ($\boldsymbol{X}^{o}-\hat{\boldsymbol{X}}$)   is orthogonal to all elements of $\mathcal{M}_{\boldsymbol{X}}$}.
@@
If on the other hand the model is nonlinear, e.g. $\boldsymbol{X}(\theta)=\sin(\boldsymbol{U}\boldsymbol{\theta})$  , then $\mathcal{M}_{\boldsymbol{X}}$} is not in general convex, much less a subspace of $\mathcal{H}$. In our example in fact ${\mathcal{M}_{\boldsymbol{X}}}:=\{\boldsymbol{X}(\boldsymbol{\theta})\}$ is not convex since for $\boldsymbol{\theta}^{1},\boldsymbol{\theta}^{2}\in\boldsymbol{\Theta},\boldsymbol{\theta}^{1}\neq\boldsymbol{\theta}^{2}$ we have $\boldsymbol{X}(\theta^{1})+\boldsymbol{X}(\theta^{2})=\sin(\boldsymbol{U}\boldsymbol{\theta}^{1})+\sin(\boldsymbol{U}\boldsymbol{\theta}^{2})\notin{\mathcal{M}_{\boldsymbol{X}}}$. Hence $\mathcal{M}_{\boldsymbol{X}}$} is not a vector space, not a subspace of the Hilbert space $\mathcal{H}=L^{2}(\Omega,\mathcal{F},P)$, no projection operator $\mathcal{P}_{{\mathcal{M}}_{\boldsymbol{X}}}$ can be defined and no unique “best” estimate $\hat{\boldsymbol{X}}$ can be obtained.
@@
However, assume $\exists$ open $\mathcal{U}\subset{\mathcal{M}_{\boldsymbol{X}}}$ and $\hat{\boldsymbol{X}}=\boldsymbol{X}(\hat{\boldsymbol{\theta}})\in\mathcal{U}$ such that $\forall\boldsymbol{X}\in\mathcal{U}$, $\boldsymbol{X}\neq\hat{\boldsymbol{X}}$ it is $\mathrm{d}({\boldsymbol{X}}^{o},\hat{\boldsymbol{X}})<\mathrm{d}(\boldsymbol{X}^{o},\boldsymbol{X})$, i.e. assume there exists an isolated point $\hat{\boldsymbol{X}}\in{\mathcal{M}_{\boldsymbol{X}}}$ locally closest to ${\boldsymbol{X}}^{o}$. Under this hypothesis, that a random variable $\hat{\boldsymbol{X}}$ exists, of minimal distance from $\boldsymbol{X}^{o}$ within a neighborhood (in the topology on $\mathcal{H}$ induced by the norm), then we could consider working on a local linear approximation to the manifold $\mathcal{M}_{\boldsymbol{X}}$} at $\hat{\boldsymbol{X}}$, that is we could consider constructing the tangent space to the manifold $\mathcal{M}_{\boldsymbol{X}}$} at that locally optimal point. To do so, we need to use the concept of the tangent space to a generic topological manifold at one of its points. Any good introductory text in differential geometry[Tu2010, Lee2009, Lang1995] will provide the relevant definitions and proofs.
A clarification is needed here: when speaking of a “tangent space” to a manifold in a differential-geometric context what is usually meant is a suitable vector space (of differential operators, see below and Appendix 6.4), which we will call in the present instance direction space and denote by $T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}$. However, we are interested here in the affine “space” of random variables $\mathcal{T}_{\hat{\boldsymbol{X}}}=\hat{\boldsymbol{X}}+T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}$  locally approximating $\mathcal{M}_{\boldsymbol{X}}$ in a neighborhood of $\hat{\boldsymbol{X}}$: we will thus call $\mathcal{T}_{\hat{\boldsymbol{X}}}$ the tangent space, even though, clearly, it does not have in general the structure of a vector space.
Now, the Hilbert projection theorem extends naturally to affine subspaces by translation, see Appendix 6.1 for details: here we will directly deal with the local approximation $\mathcal{T}_{\hat{\boldsymbol{X}}}$ to $\mathcal{M}_{\boldsymbol{X}}$} at $\hat{\boldsymbol{X}}$, calling it, with a slight abuse of language, a “subspace” of $\mathcal{H}$, see Fig.6.
@@
We can build the direction space $T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}$ by considering the set of velocities, i.e. of differential operators, which map functions in $C^{\infty}({\mathcal{M}_{\boldsymbol{X}}})$ , $f:{\mathcal{M}_{\boldsymbol{X}}}\rightarrow\mathbb{R}$,  to their directional derivatives in the direction of curves $\gamma:\mathbb{R}\rightarrow{\mathcal{M}_{\boldsymbol{X}}}$, i.e.:
$T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}=\{\vartheta_{\gamma,\hat{\boldsymbol{X}}}$ such that $\hat{\boldsymbol{X}}=\gamma(\lambda)\in\mathcal{U}\subset{\mathcal{M}_{\boldsymbol{X}}}$ and $\forall f\in C^{\infty}({\mathcal{M}_{\boldsymbol{X}}})$ it is $\vartheta_{\gamma,\hat{\boldsymbol{X}}}f=(f\circ\gamma)^{\prime}|_{\lambda}\}$.
Notice that here the curve $\gamma$ is just one representative of a family of equivalent curves $\{\gamma^{\bullet}:\gamma^{\bullet}(\lambda^{\bullet})=\gamma(\lambda)=\hat{\boldsymbol{X}},(f\circ\gamma^{\bullet})^{\prime}|_{\lambda^{\bullet}}=(f\circ\gamma)^{\prime}|_{\lambda}\}$.
It is now necessary both to establish a basis for $T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}$ and to clarify the correspondence between the differential operators in $T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}$ and the random variables that ought to belong to $\mathcal{T}_{\hat{\boldsymbol{X}}}$.
We can do this using the “canonical” chart from the manifold $\mathcal{M}_{\boldsymbol{X}}$} to $\mathbb{R}^{q}$ induced by the parametrization of the model. We can in general define a chart-induced basis of a tangent space: in the present case we will compute the basis for $T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}$ induced by the canonical chart $\boldsymbol{X}^{-1}$, similarly to the canonical chart map from $\mathbb{R}^{n}$ to $\mathbb{R}^{q}$ already introduced at line LABEL:canonicalchartmap1 of page LABEL:canonicalchartmap1:
$\mathcal{M}_{\boldsymbol{X}}$$\mathbb{R}^{q}$$\boldsymbol{X}^{-1}$$\boldsymbol{X}$
@@
Clearly we must require sufficient regularity in the model map $\boldsymbol{X}$ to do this, such as being continuously differentiable with a continuously differentiable inverse. Notice that $\boldsymbol{X}^{-1}$ is a chart, indeed a global chart, mapping the whole of $\mathcal{M}_{\boldsymbol{X}}$} to $\mathbb{R}^{q}$, since $\mathcal{M}_{\boldsymbol{X}}$} is generated by $\boldsymbol{X}(\boldsymbol{\Theta})$, with $\boldsymbol{\Theta}\in\mathbb{R}^{q}$, because by definition for every admissible $\boldsymbol{\theta}$ there is a point $\boldsymbol{X}(\boldsymbol{\theta})\in{\mathcal{M}_{\boldsymbol{X}}}$.
We can also say, therefore that:
@@
Summarizing what is more precisely detailed in Appendix 6.4, we can thus use as basis of the tangent space $T_{\hat{\boldsymbol{X}}}{{\mathcal{M}_{\boldsymbol{X}}}}$ the basis induced by the chart map $(\boldsymbol{X}^{-1})$:
$\{(\cdot\circ((\boldsymbol{X}^{-1})^{-1})^{j})^{\prime}|_{\hat{\boldsymbol{X}}}\}_{j=1,...,q}=\{(\cdot\circ\boldsymbol{X}^{j})^{\prime}|_{\hat{\boldsymbol{X}}}\}_{j=1,...,q}\,\,.$